# Exotic Doublon-Holon Pairing State in Photodoped Mott Insulators

### v1

@@ Abstract
To examine the pairing symmetry, following studies of chemically doped systems [57, 63, 64, 65, 66, 67], we introduce $\hat{\Delta}^{\mathrm{c}}_{j}=\sum_{\sigma,\beta=0,1}(-1)^{\beta}\left(1-\hat{n}_{j+\bar{\beta},0;\bar{\sigma}}\right)\hat{c}^{\dagger}_{j+\bar{\beta},0;\sigma}\hat{c}_{j+\beta,0;\sigma}\hat{n}_{j+\beta,0;\bar{\sigma}}/2$ for the DH pair formation along the chain direction and calculate the rung-chain DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$. As shown in Fig. 2(b), the rung-rung correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and the rung-chain correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibit opposite signs. Similar sign inversion between $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ appears in the SC correlations in chemically doped ladders, which was taken as a signature of $d$-wave pairing [57]. This intriguing finding suggests that the DH pairing correlation in the 2D square lattice would also have the $d$-wave symmetry and thus potentially bears a close similarity to the equilibrium SC state in chemically doped MIs.
In other regions of the $(n_{d},V)$ phase diagram, various phases can dominate. These include the CDW and $\eta$-pairing nonequilibrium phases, which have been previously studied in the 1D chain [43, 44]. Our results confirm their presence in the ladder geometry, see Fig. 1(a), and also reveal a spin-singlet phase distinct from the SDW phase in the 1D chain. As shown in Fig. 1(a), for $t_{\perp}=t_{\parallel}$, we observe the spin-singlet phase at $n_{d}=0$, where $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(s)}_{J}$. In this phase, the spin correlations decay exponentially [86], implying the presence of the spin gap due to the rung spin-singlet formation. Similarly to the 1D chain [43, 44], the staggered on-site pairing, i.e., $\eta$-pairing, correlations become dominant in the small $V$ region, and the CDW correlations become dominant in the large $n_{d}$ and large $V$ region [86]. The $\eta$-pairing and CDW are degenerate at $n_{d}=0.5$ and $V=0.2t_{\parallel}$. Since the system at $n_{d}=0.5$ is composed of doublons and holons without singly occupied sites, the effective model $\hat{\mathcal{H}}$ reduces to ^$\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}$, which is equivalent to the XXZ model. The point at which the $\eta$-pairing and CDW are degenerate corresponds to the transition point between the XY and Ising phases in the XXZ model [43].
@@
Effective model— Here, we explain the origin of the phase factor $(-1)^{r}$ in the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$. Based on the above results, we can conclude that an important aspect of DH-pairing is the formation of DH $\eta$-triplets on the spin-singlet background. Thus, we discuss these behaviors in a minimal model with local state space restricted to rung spin-singlet and rung DH $\eta$-triplet. We derive a minimal model from the local rung approximation [92], ^$\hat{\mathcal{H}}_{\rm LRA}=\left[\hat{\mathcal{H}}^{(0)}_{t_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}\right]+\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$, considering the strong rung coupling regime with the interrung hopping ^$\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}$ as the perturbation. The second-order perturbation theory gives the effective Hamiltonian for the two-level system
^$\hat{\mathcal{H}}_{\rm min}=K\sum_{j}\left(\hat{\mathbf{\Delta}}_{j}\cdot\hat{\mathbf{\Delta}}_{j+1}-\frac{1}{4}\right),$ (5)
where $K=2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}+2t_{\perp})+2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}-2t_{\perp})$, see Supplemental Material for details [86]. The rung DH $\eta$-triplet and rung spin-singlet states define the effective up and down pseudospin states, respectively, see inset of Fig. 4(b). $\hat{\mathbf{\Delta}}_{j}$ is a vector composed of the pseudospin operators $\hat{\Delta}^{+}_{j}$, $\hat{\Delta}^{-}_{j}$, and $\hat{\Delta}^{z}_{j}$. As before, $\hat{\Delta}^{+}_{j}=\hat{\Delta}^{r\dagger}_{j}$ maps the rung spin-singlet state to the DH $\eta$-triplet state. The fixed number of doublons in photodoped MIs corresponds to the fixed pseudo magnetization in Eq. (5).
Figure 4: (a) Linear scale plot and (b) log-log plot of the rung-rung DH pairing correlation function with varying $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$. Here, $t_{\perp}$ is used as the unit of energy. We set $n_{d}=0.05$, $J_{\perp}=0.4t_{\perp}$, and $V_{\perp}=1.8t_{\perp}$. The interrung parameters are varied as $t_{\parallel}=(0.1+0.9p)t_{\perp}$, $J_{\parallel}=0.4pt_{\perp}$, and $V_{\parallel}=1.8pt_{\perp}$, where $0\leq p\leq 1$. Inset: correspondence between the photodoped ladder and the 1D antiferromagnetic Heisenberg model. The blue and purple ellipses represent the rung spin-singlet and rung DH $\eta$-triplet, respectively, as shown in Fig. 1(b).
This effective model assumes $V_{\perp}+J_{\perp}>|2t_{\perp}|$ to set the spin-singlet and $\eta$-triplet states to the lowest-energy configurations in the independent rung limit [86]. Under this condition, $K>0$, and therefore Eq. (5) is equivalent to the 1D antiferromagnetic Heisenberg model at a certain magnetization. The lowest-energy state of Eq. (5) coincides with the ground state of the Heisenberg model under a magnetic field. Due to the equivalence with the antiferromagnetic Heisenberg model, $\braket{\hat{\Delta}^{+}_{j+r}\hat{\Delta}^{-}_{j}}$ in Eq. (5) must have the phase factor $(-1)^{r}$, explaining the sign alternation of the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the photodoped ladder.
Going beyond the above analytical perturbative arguments, we study the stability of the sign alternation in $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{i}\hat{\Delta}^{\mathrm{r}}_{j}}$ as a function of the $t_{\parallel},J_{\parallel},V_{\parallel}$ parameters of the full model $\hat{\mathcal{H}}$ numerically. Figure 4 shows the DH pairing correlations as $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$ increase linearly, departing from ^$\hat{\mathcal{H}}_{\rm LRA}$. As expected from the effective two-level model in Eq. (5), the DH pairing correlation is well developed when different rungs are only weakly connected ($p=0$). Although the absolute value of the correlation gradually decreases at larger $p$ as the parameters approach the original $\hat{\mathcal{H}}$, the sign alternation in the correlation function remains, confirming that ^$\hat{\mathcal{H}}_{\mathrm{min}}$ captures the essential features of the pairing correlations observed in the DH pairing phase.
[…]
@@
In contrast to the $\eta$-pairing phase that prominently appears in the $V\sim 0$ and large $n_{d}$ region, the DH pairing state can appear in the large $V$ and small $n_{d}$ region. Because strong photoexcitation to achieve high $n_{d}$ is not required and $V$ is nonzero in real materials, the DH pairing state is more experimentally accessible. For instance, ladder-type cuprates [93, 94, 95, 96] may serve as hosts for the DH pairing state. While the implementation of nearest-neighbor Coulomb interaction is a challenging ingredient for the realization with cold atoms, the first step towards this has been achieved for the Bose-Hubbard model [97].
@@
Although photodoping is inherently a nonequilibrium phenomenon, we have approximately mapped a photodoped state to the lowest-energy state in the pseudoequilibrium condition. To further validate the emergence of the DH pairing phase and its lifetime, it would be necessary to demonstrate the time-dependent behavior of the DH pairing correlations in the optically driven Hubbard ladder using time-evolution methods [98, 99, 100, 101, 102]. These investigations remain as future challenges.
Acknowledgements.
@@
We thank S. Ejima, Y. Murakami, S. Nishimoto, and T. Sato for fruitful discussions. This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grants No. JP20H01849, No. JP24K06939, No. JP24H00191 (T.K.), No. JP22K04907 (K.K.), and No. JP24K01333. R.U. was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program”. M.S. acknowledges the support from the UK EPSRC award under the Agreement EP/Y005090/1. Z.L. acknowledges the support by the QuantERA II JTC 2021 grant T-NiSQ by MVZI, the P1-0044 program of the Slovenian Research Agency, and ERC StG 2022 project DrumS, Grant Agreement 101077265. D.G. is supported by the Slovenian Research and Innovation Agency (ARIS) under Programs No. P1-0044, No. J1-2455, and No. MN-0016-106. Our calculations were performed using the ITensor library [103, 104].

### v2

@@ Abstract
Figure 2 shows the correlation functions at $n_{d}=0.2$ and $V=0.2t_{\parallel}$ in the isotropic ladder ($t_{\perp}=t_{\parallel}$, $V_{\perp}=V_{\parallel}=V$). The log-log plot in Fig. 2(a) shows that the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibits a power-law decay, while the spin ($\hat{O}_{j}=\hat{n}_{j,\alpha;\uparrow}-\hat{n}_{j,\alpha;\downarrow}$), charge ($\hat{O}_{j}=\hat{n}_{j,\alpha}-1$), and on-site pair ($\hat{O}_{j}=\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$) correlations [96] decay exponentially, indicating that the DH pairing correlation is dominant. As shown in Fig. 2(b), the rung-rung correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and the rung-chain correlation $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ exhibit opposite signs. Similar sign inversion between superconducting rung-rung and rung-chain correlators appears in chemically doped ladders [91], which was taken as a signature of $d$-wave pairing [61]. This intriguing finding suggests that the DH pairing correlation in 2D systems potentially bears a close similarity to the equilibrium $d$-wave SC state in chemically doped MIs.
@@
The emergence of long-range correlations along the chain is a nontrivial result that cannot be explained by local pair formation alone. Our DMRG calculations demonstrate that the DH pairing correlations can grow over long distances, revealing not only local exciton formation but also a long-range development of the pairing correlation. The DH pairing state is stable in a wide regime of doublon density $n_{d}$ and Coulomb interaction $V$ as shown in Fig. 1(a) [97]. In other regions of the $(n_{d},V)$ phase diagram, various phases can dominate. Their detailed properties are presented in the Supplemental Material [91]. Similar to the 1D chain [40, 41], the staggered on-site pairing, i.e., $\eta$-pairing, correlations become dominant in the small $V$ region, and the CDW correlations become dominant in the large $n_{d}$ and large $V$ region. At $n_{d}=0$, we observe the spin-singlet phase, which is distinct from the SDW phase in the 1D chain, because a coupled spin chain with a spin gap [98] is stabilized.
Figure 3: (a) Rung-rung DH pairing correlation function for various $t_{\perp}/t_{\parallel}$ values at $L=120$, $n_{d}=0.1$ and $V=V_{\perp}=V_{\parallel}=0.4t_{\parallel}$, where $J_{\perp}/J_{\parallel}=(t_{\perp}/t_{\parallel})^{2}$ with $J_{\parallel}=0.4t_{\parallel}$. (b) Linear scale and (c) log-log plots of the DH pairing correlation functions in the four-leg cylinder, where $n_{d}=0.125$, $J=0.4t_{\rm h}$, and $V=0.8t_{\rm h}$.
To examine the evolution of pairing tendency from 1D to ladder geometry, Fig. 3(a) shows the DH pairing correlations for various $t_{\perp}/t_{\parallel}$ values. The DH pairing correlations increase as $t_{\perp}/t_{\parallel}$ increases. While the SDW correlation is dominant in the small-$t_{\perp}$ regime [91], the DH pairing correlation becomes dominant at $t_{\perp}/t_{\parallel}\gtrsim 1$. This $t_{\perp}$ dependence suggests that DH pairing does not arise from one-dimensionality. To extrapolate our observation towards 2D systems, we compute the DH pairing correlation on a cylinder square lattice with $8\times 4$ sites and periodic boundary conditions in the $y$ direction. Figures 3(b) and 3(c) present the DH pairing correlations calculated by the DMRG method. Results on a cylinder are consistent with those on a ladder: we find a slow power-law decay of the pairing correlation, accompanied by the sign alternations. This suggests that DH pairing should not be limited to the ladder system and should persist in 2D.
Effective model. Having observed a stable staggered nature of the DH pairing, we derive here a minimal model that can explain the phase factor of $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the ladder setup. The $t_{\perp}/t_{\parallel}$ dependence shown in Fig. 3(a) suggests that a background of rung spin singlets, stabilized by the strong interchain coupling via $t_{\perp}$, is favorable for the emergence of the DH pairing phase on top of it. The preferable formation of rung spin singlets as $t_{\perp}/t_{\parallel}$ increases is confirmed in the Supplemental Material [91]. Since the formation of DH $\eta$ triplets on the spin-singlet background is important for DH pairing characterized by the $\hat{\Delta}^{r}_{j}$ operator [see Fig. 1(d)], we discuss these behaviors in an effective model with local state space restricted to rung spin singlet and rung DH $\eta$ triplet.
We derive a minimal model from the local rung approximation [99], $\hat{\mathcal{H}}_{\rm LRA}=\hat{\mathcal{H}}_{\perp}+\hat{\mathcal{T}}_{\parallel}$, considering the strong rung coupling in $\hat{\mathcal{H}}_{\perp}$ [rung-direction component in Eq. (Doublon-holon pairing state in photodoped Mott insulators)] with the interrung hopping $\hat{\mathcal{T}}_{\parallel}$ as the perturbation. The second-order perturbation theory gives the effective Hamiltonian for the two-level system
$\hat{\mathcal{H}}_{\rm min}=K\sum_{j}\left(\hat{\mathbf{\Delta}}_{j}\cdot\hat{\mathbf{\Delta}}_{j+1}-\frac{1}{4}\right),$ (5)
where $K=2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}+2t_{\perp})+2t_{\parallel}^{2}/(V_{\perp}+J_{\perp}-2t_{\perp})$ (see Supplemental Material [91] for details). The rung DH $\eta$-triplet and rung spin-singlet states define the effective up and down pseudospin states, respectively [see the inset of Fig. 4(b)]. $\hat{\mathbf{\Delta}}_{j}$ is a vector composed of the pseudospin operators $\hat{\Delta}^{+}_{j}$, $\hat{\Delta}^{-}_{j}$, and $\hat{\Delta}^{z}_{j}$. As before, $\hat{\Delta}^{+}_{j}=\hat{\Delta}^{r\dagger}_{j}$ maps the rung spin-singlet state to the DH $\eta$-triplet state. The fixed number of doublons in a photodoped MI corresponds to the fixed pseudomagnetization in Eq. (5). Since this effective model is derived under the condition $K>0$ [91], Eq. (5) is equivalent to the 1D antiferromagnetic Heisenberg model at a certain magnetization. The lowest-energy state of Eq. (5) coincides with the ground state of the Heisenberg model under a magnetic field. Due to the equivalence with the antiferromagnetic Heisenberg model, $\braket{\hat{\Delta}^{+}_{j+r}\hat{\Delta}^{-}_{j}}$ in Eq. (5) must have the phase factor $(-1)^{r}$, explaining the sign alternation of the DH pairing correlation $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j+r}\hat{\Delta}^{\mathrm{r}}_{j}}$ in the photodoped ladder.
Figure 4: (a) Linear scale plot and (b) log-log plot of the rung-rung DH pairing correlation function with varying $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$. Here, $t_{\perp}$ is used as the unit of energy. We set $n_{d}=0.05$, $J_{\perp}=0.4t_{\perp}$, and $V_{\perp}=1.8t_{\perp}$. The interrung parameters are varied as $t_{\parallel}=(0.1+0.9p)t_{\perp}$, $J_{\parallel}=0.4pt_{\perp}$, and $V_{\parallel}=1.8pt_{\perp}$, where $0\leq p\leq 1$. Inset: Correspondence between the photodoped ladder and the 1D antiferromagnetic Heisenberg model. The blue and purple symbols represent the rung spin singlet and rung DH $\eta$ triplet, respectively, as shown in Fig. 1(b).
Going beyond the above analytical perturbative arguments, we study the stability of the sign alternation in $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{i}\hat{\Delta}^{\mathrm{r}}_{j}}$ as a function of the $t_{\parallel},J_{\parallel},V_{\parallel}$ parameters of the full model $\hat{\mathcal{H}}$ numerically. Figure 4 shows the DH pairing correlations as $t_{\parallel}$, $J_{\parallel}$, and $V_{\parallel}$ increase linearly, departing from $\hat{\mathcal{H}}_{\rm LRA}$. As expected from the effective two-level model in Eq. (5), the DH pairing correlation is well developed when different rungs are only weakly connected ($p=0$). Although the absolute value of the correlation gradually decreases at larger $p$ as the parameters approach the original $\hat{\mathcal{H}}$, the sign alternation in the correlation function remains, confirming that $\hat{\mathcal{H}}_{\mathrm{min}}$ captures the essential features of the pairing correlations observed in the DH pairing phase.
@@
The DH pairing state can appear in the small $n_{d}$ region, thus strong photoexcitation to achieve high $n_{d}$ is not required, and heating effects should be less destructive. The Coulomb interaction $V$ cooperatively contributes to the emergence of the DH pairing state; such an interaction is natively present in many correlated materials. For instance, ladder-type cuprates [100, 101, 102, 103] may serve as hosts for the DH pairing state. Having charted the DH pairing in a 2D setting, various other materials, such as cuprates with the 2D square structure, hold the potential to exhibit DH pairing. While the implementation of a nearest-neighbor Coulomb interaction is a challenging ingredient for the realization with cold atoms, the first step towards this has been achieved for the Bose-Hubbard model [104].
@@
Although photodoping is inherently a nonequilibrium phenomenon, we have approximately mapped a photodoped state to the lowest-energy state in the pseudoequilibrium condition. To further validate the emergence of the DH pairing phase and its lifetime, it would be necessary to demonstrate the time-dependent behavior of the DH pairing correlations in the optically driven Hubbard ladder using time-evolution methods [105, 106, 107, 108, 109]. Having obtained a promising behavior on a small 2D cylinder, further studies of the $d$-wave nature of pairing should be performed on larger 2D lattices. These investigations remain as future challenges.
@@
Acknowledgments. We thank S. Ejima, Y. Murakami, S. Nishimoto, K. Pradhan, and T. Sato for fruitful discussions. This work was supported by Grants-in-Aid for Scientific Research from JSPS, KAKENHI Grants No. JP20H01849 (T.K.), No. JP24K06939 (T.K.), No. JP24H00191 (T.K.), No. JP22K04907 (K.K.), and No. JP24K01333. R.U. was supported by the Program for Leading Graduate Schools: “Interactive Materials Science Cadet Program” and JST SPRING Grant No. JPMJSP2138. M.S. acknowledges the support from the U.K. EPSRC award under Agreement No. EP/Y005090/1. Z.L. acknowledges the support by the QuantERA II JTC 2021 Grant T-NiSQ by MVZI, the P1-0044 program of the Slovenian Research Agency, and ERC StG 2022 project DrumS, Grant Agreement No. 101077265. D.G. is supported by the Slovenian Research and Innovation Agency (ARIS) under Programs No. P1-0044, No. J1-2455, and No. MN-0016-106. Our calculations were performed using the ITensor library [110, 111].
Data availability. The data that support the findings of this article are openly available [112].